28/39 divided by 20/9
Answer by amoresroy(361)

28/39 divided by 20/9
(28/39) / (20/9) = (28/39) * (9/20)
= (28*9) / (39*20)
= (4*7*3*3) / (3*13*4*5)
4*3 cancels out
= (7*3)/ (13*5)
= 21/65